OpenStudy (anonymous):

if p= 3i+2j- 4k, what is |P|

OpenStudy (amistre64):

it is the length of the vector

OpenStudy (pdoctor):

Is |P| = Magnitude or absolute value?

OpenStudy (amistre64):

squre then add them and sqrt them

OpenStudy (amistre64):

sqrt(29) if i see it right

OpenStudy (turingtest):

\[\left| P \right|=\sqrt{3^2+2^2+(-4)^2}=5.385\]

OpenStudy (turingtest):

magnitude

OpenStudy (turingtest):

yes sqr(29) is right

OpenStudy (amistre64):

yay!! now i can be wrong for the rest of the day ;)

OpenStudy (turingtest):

lol..